OpenStudy (anonymous):

I am confused what does one to one mean. ex. If the function is one to one find the inverse. g(x)= square root( x+6),x is greater than our equal to -6

OpenStudy (anonymous):

\[g(x)=\sqrt{x+6}\] and you want the inverse put \[x=\sqrt{y+6}\] and solve for \(y\)

OpenStudy (anonymous):

you get \[x^2=y+6\]\[x^2-6=y\] so \[g^{-1}(x)=x^2-6\]

OpenStudy (anonymous):

function is one to one means a) it passes the "horizontal line test" or b) if \(f(a)=f(b)\) then \(a=b\) in other words no two different inputs can give the same output